# Double-coset zeta functions for groups acting on trees

### v1

@@ 6. The reciprocal of $\mathcal{Z}_{\Gamma,u\to u}(s)$
In view of Section 7.2, we present some formulae involving the reciprocal of the function $\mathcal{Z}_{\Gamma,u\to u}(s)$, for $u\in\Gamma$, introduced in Definition 5.9. Recall that this function is a generalisation to weighted graphs of $\zeta_{G,G_{t},G_{t}}(s)$, where $(G,T)$ is a weakly locally $\infty$-transitive group action on a locally finite tree and $t\in T$ (cf. Remark 5.9.2).
@@
Let $\Gamma$ be a non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Consider $\mathcal{E}(s)$, for $s\in\mathbb{C}$, and $\{e_{u}\}_{u\in\Gamma}$ as in Definition 5.10. For all $c\in V\Gamma$ and $a\in E\Gamma$, define
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$], and denote by $I$ the identity matrix on $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$. Then, for all $c\in V\Gamma$ and $s\in\mathbb{C}$ such that $I-\mathcal{G}_{c}(s)$ is invertible, we have
@@
The statements now follow from Fact 5.13. ∎
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider two subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$, for some $c\in V\Gamma$. Then,
@@
(6.4) $\mathcal{Z}_{\Gamma_{i},c\to c}(s)^{-1}=1-\sum_{a\in o^{-1}(c)\cap E\Gamma_{i},\atop b\in t^{-1}(c)\cap E\Gamma_{i}}\omega(a)^{-s}(I_{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))^{-1}(a,b).$
Moreover, for all $a,b\in E\Gamma$ with $t(a)=c=o(b)$ we have $(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=\mathbbm{1}_{\{a\}}(b)-(\omega(b)-\mathbbm{1}_{\{\bar{a}\}}(b))^{-s}+\omega(b)^{-s}$ and then
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{c\}$, for some vertex $c\in V\Gamma$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
@@
The hypotheses of Proposition 6.4 are satisfied for every $c\in V\Gamma$ if, for instance, $\Gamma$ is a connected graph without $n$-cycles, for every $n\geq 2$.
@@
Let $\{\Xi_{i}\}_{i\in I}$ be the collection of all connected components of the graph $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})$. Recall that $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})=\bigcup_{i\in I}\Xi_{i}$ and the union is disjoint. For every $i\in I$, there is exactly one edge $a_{i}\in o^{-1}(c)$ such that $t(a_{i})\in\Xi_{i}$. In fact, assume that there are $a,b\in o^{-1}(c)$ with $a\neq b$ and $x=t(a),y=t(b)\in V\Xi_{i}$. Then the reduced path $[x,y]$ as in Remark 2.1.1 is contained in $\Xi_{i}$. Therefore, $a\cdot[x,y]\cdot\bar{b}$ is a cycle of length $\geq 2$ in $\Gamma$, impossible.
Consider subsets $\mathcal{E}_{1},\mathcal{E}_{2}$ of $o^{-1}(c)$ with the following properties: $\mathcal{E}_{1}\cap\mathcal{E}_{2}=\emptyset$, $o^{-1}(c)=\mathcal{E}_{1}\cup\mathcal{E}_{2}$ and, for every $k\in\{1,2\}$, every $1$-loop $a$ in $\mathcal{E}_{k}$ satisfies that $\bar{a}\in\mathcal{E}_{k}$. Provided ¯$\overline{\mathcal{E}}_{k}=\{\bar{a}\mid a\in\mathcal{E}_{k}\}$ for every $k\in\{1,2\}$, note that $(\mathcal{E}_{1}\cup\overline{\mathcal{E}}_{1})\cap(\mathcal{E}_{2}\cup\overline{\mathcal{E}}_{2})=\emptyset$. Moreover, if $|o^{-1}(c)|\geq 2$, take $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ so that $\mathcal{E}_{1}\neq\emptyset$ and $\mathcal{E}_{2}\neq\emptyset$. For $k\in\{1,2\}$, set also $I_{k}:=\{i\in I\mid a_{i}\in\mathcal{E}_{k}\}$. Note that $I_{1}\cap I_{2}=\emptyset$ and $I=I_{1}\cup I_{2}$. For every $k\in\{1,2\}$, define the following subgraph of $\Gamma$:
$\Gamma_{k}:=\{c\}\cup\mathcal{E}_{k}\cup\overline{\mathcal{E}}_{k}\cup\bigcup_{i\in I_{k}}\Xi_{i}.$
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{3}:=\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Then,
@@
Note that $c\in V\Lambda_{1}$, $d\in V\Lambda_{2}$ and $V\Lambda_{1}\cap V\Lambda_{2}=\emptyset$. In particular, no edges of $\Lambda_{1}$ end in a vertex of $\Lambda_{2}$. We claim that
@@
From (6.18) we observe that $A(a,a)=1+\alpha^{-s}$ and $B(\bar{a},\bar{a})=1+\beta^{-s}$. Moreover, by (6.18) and (6.19), we have the following decompositions in diagonal blocks:
(6.20) $\begin{split}I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s)=\begin{bmatrix}A&0\\ 0&B\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma}-\mathcal{G}_{a}^{\Gamma}(s))^{-1}=\begin{bmatrix}A^{-1}&0\\ 0&B^{-1}\end{bmatrix};\\ I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s)=\begin{bmatrix}A&0\\ 0&B(\bar{a},\bar{a})\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma_{1}}-\mathcal{G}_{a}^{\Gamma_{1}}(s))^{-1}=\begin{bmatrix}A^{-1}&0\\ 0&B(\bar{a},\bar{a})^{-1}\end{bmatrix};\\ I^{\Gamma_{2}}-\mathcal{G}_{a}^{\Gamma_{2}}(s)=\begin{bmatrix}A(a,a)&0\\ 0&B\end{bmatrix}&\quad\text{and}\quad(I^{\Gamma_{2}}-\mathcal{G}_{a}^{\Gamma_{2}}(s))^{-1}=\begin{bmatrix}A(a,a)^{-1}&0\\ 0&B^{-1}\end{bmatrix}.\end{split}$
@@
(6.23) $\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}=1-\beta^{-s}B(\bar{a},\bar{a})^{-1}-\alpha^{-s}A^{-1}(a,a)-\sum_{b\in o^{-1}(c)\cap E\Lambda_{1}}\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},b)\cdot A^{-1}(b,a).$
@@
(6.24) $\mathcal{Z}_{\Gamma_{2},a\to a}(s)^{-1}=1-\alpha^{-s}A(a,a)^{-1}-\beta^{-s}B^{-1}(\bar{a},\bar{a})-\sum_{b\in o^{-1}(d)\cap E\Lambda_{2}}\mathcal{E}^{\Gamma_{2}}(s)(a,b)\cdot B^{-1}(b,\bar{a}).$
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ satisfying $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and such that $\Lambda_{1}\cap\Lambda_{2}$ is a $1$-segment graph with edge set $\{a,\bar{a}\}$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ as in the statement. For simplicity, set $\Gamma_{3}=\Gamma_{1}\cap\Gamma_{2}$ and $\Gamma_{a}=\Lambda_{1}\cap\Lambda_{2}$. Since $\Lambda_{1}\subseteq\Gamma_{1}$ and $\Lambda_{2}\subseteq\Gamma_{2}$, we have $\Lambda_{1}\cap\Gamma_{2}=\Lambda_{1}\cap\Gamma_{3}$ and $\Lambda_{2}\cap\Gamma_{1}=\Lambda_{2}\cap\Gamma_{3}$. Therefore, the following decompositions hold:
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$]. Consider $a\in E\Gamma$ with $o(a)=:c\neq d:=t(a)$, and assume that $c$ is a terminal vertex in $\Gamma$. Put $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and denote by $\Lambda$ the graph obtained from $\Gamma$ by removing $a$, $\bar{a}$ and $c$. Then,
@@
Let $s\in\mathbb{C}$ such that $\mathcal{Z}_{\Gamma,a\to a}(s)\neq 0$, and consider a total order $\leq$ on $E\Gamma$ such that $a<\bar{a}<b$ for all $b\in E\Lambda$. Then $\mathcal{E}^{\Gamma}(s)$ admits the following block decomposition:
@@
where $\underline{0}$ denotes the row zero vector in $\mathbb{C}^{|E\Lambda|}$. Moreover, denote by $E_{21}$ and $\underline{1}\in\mathrm{Mat}_{2}(\mathbb{C})$ the elementary matrix associated to $(2,1)$ and the matrix with all the entries equal to $1$, respectively. Hence,
$\begin{split}\mathcal{U}_{cc}^{\Gamma}(s)&=(\alpha+1)^{-1}(e_{\bar{a}}^{\Gamma})^{t}(e_{a}^{\Gamma})=\begin{bmatrix}(\alpha+1)^{-s}E_{21}&0_{2\times|E\Lambda|}\\ 0_{|E\Lambda|\times 2}&0_{|E\Lambda|\times|E\Lambda|}\end{bmatrix};\\ \mathcal{U}_{aa}^{\Gamma}(s)&=(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})^{t}(e_{a}^{\Gamma}+e_{\bar{a}}^{\Gamma})\mathcal{E}^{\Gamma}(s)=\begin{bmatrix}\underline{1}\cdot A&\underline{1}\cdot B\\ 0_{|E\Lambda|\times 2}&0_{|E\Lambda|\times|E\Lambda|}\end{bmatrix}.\end{split}$
@@
(6.34) $\begin{split}I^{\Gamma}-\mathcal{E}^{\Gamma}(s)&=\begin{bmatrix}I_{2}-A&-B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I^{\Gamma}-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{c,c}^{\Gamma}(s)&=\begin{bmatrix}I_{2}-A+(\alpha+1)^{-s}E_{21}&-B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I^{\Gamma}-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{a,a}^{\Gamma}(s)&=\begin{bmatrix}I_{2}+(\underline{1}-I_{2})A&(\underline{1}-I_{2})B\\ -C&I_{\Lambda}-D\end{bmatrix};\\ I_{2}-A=\begin{bmatrix}1&-\beta^{-s}\\ -\alpha^{-s}&1\end{bmatrix};&\qquad I_{2}+(\underline{1}-I_{2})A=\begin{bmatrix}1+\alpha^{-s}&0\\ 0&1+\beta^{-s}\end{bmatrix}.\end{split}$
@@
(6.35) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}&=\frac{\det(I^{\Gamma}-\mathcal{E}^{\Gamma}(s))}{\det(I-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{c,c}^{\Gamma}(s))}\\ &=\frac{\det\Big{(}I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=:X}\Big{)}}{\det\Big{(}I_{2}-A+(\alpha+1)^{-s}E_{21}-\underbrace{B(I_{\Lambda}-D)^{-1}C}_{=X}\Big{)}}\\ &=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1-X(a,a))(1-X(\bar{a},\bar{a}))+(\xi(\alpha,s)-X(\bar{a},a))(\beta^{-s}+X(a,\bar{a}))}.\end{split}$
@@
(6.36) $\begin{split}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=\frac{\det(I^{\Gamma}-D)}{\det(I-D+\mathcal{U}_{a,a}^{\Gamma}(s))}\\ &=\frac{\det\Big{(}I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=X}\Big{)}}{\det\Big{(}I_{2}+(\underline{1}-I_{2})A+\underbrace{(\underline{1}-I_{2})B(I_{\Lambda}-D)^{-1}C}_{=:Y}\Big{)}}\\ &=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1+\alpha^{-s}+Y(a,a))(1+\beta^{-s}+Y(\bar{a},\bar{a}))-Y(a,\bar{a})Y(\bar{a},a)}.\end{split}$
@@
Moreover, by Theorem 5.12, Fact 5.13 and since $\mathcal{E}^{\Lambda}(s)=D$,
$\begin{split}X(a,\bar{a})=Y(\bar{a},\bar{a})&=B_{1}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{2}\\ &=(\beta+1)^{-s}\cdot\sum_{b\in o^{-1}(d)\cap E\Lambda}\omega(b)^{-s}e_{b}^{\Lambda}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}(e_{d}^{\Lambda})^{t}\\ &=(\beta+1)^{-s}(\mathcal{Z}_{\Lambda,d\to d}(s)-1).\qed\end{split}$
@@
Let $(\Gamma,\omega)$ satisfy Setting [$\Gamma$], and consider $a\in E\Gamma$ with $o(a)=t(a)=c$. Set $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and consider the subgraph of $\Gamma$ given by $\Lambda:=\Gamma\setminus\{a,\bar{a}\}$. Then,
@@
The picture below sketches the setting of Proposition 6.10. The edges of $\Lambda$ are dashed.
@@
(6.38) $\begin{split}I_{2}-A&=\begin{bmatrix}1-(\alpha+1)^{-s}&-\beta^{-s}\\ -\alpha^{-s}&1-(\beta+1)^{-s}\end{bmatrix};\\ I_{2}+(\underline{1}-I_{2})A&=\begin{bmatrix}1+\alpha^{-s}&(\beta+1)^{-s}\\ (\alpha+1)^{-s}&1+\beta^{-s}\end{bmatrix}.\end{split}$
@@
In this case $X=Y$, because $B_{1}=B_{2}$ and then $(\underline{1}-I_{2})B=B$. Moreover, recalling Theorem 5.12 and Fact 5.13, we have

### v2

@@ 6. The reciprocal of $\mathcal{Z}_{\Gamma,u\to u}(s)$
In view of Section 7.2, we present some formulae involving the reciprocal of the function $\mathcal{Z}_{\Gamma,u\to u}(s)$, for $u\in\Gamma$, introduced in Definition 5.11. Recall that this function is a generalisation of $\zeta_{G,G_{t},G_{t}}(s)$, where $(G,T)$ is a weakly locally $\infty$-transitive group action on a locally finite tree and $t\in T$ (cf. Remark 5.11.2).
@@
Let $\Gamma$ be a non-empty graph with an edge weight $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$. Consider $\mathcal{E}(s)$, for $s\in\mathbb{C}$, and $(e_{u})_{u\in\Gamma}$ as in Definition 5.12. For all $c\in V\Gamma$ and $a\in E\Gamma$, define
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$], and denote by $I$ the identity matrix on $\mathrm{Mat}_{|E\Gamma|}(\mathbb{C})$. Then, for all $c\in V\Gamma$ and $s\in\mathbb{C}$ such that $I-\mathcal{G}_{c}(s)$ is invertible, we have
@@
The statements now follow from Fact 5.15. ∎
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$]. Consider two subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}=\{c\}$, for some $c\in V\Gamma$. Then,
@@
(6.4) $\mathcal{Z}_{\Gamma_{i},c\to c}(s)^{-1}=1-\sum_{a\in o^{-1}(c)\cap E\Gamma_{i},\atop b\in t^{-1}(c)\cap E\Gamma_{i}}\omega(a)^{-s}(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))^{-1}(a,b).$
Moreover, for all $a,b\in E\Gamma$ with $t(a)=c=o(b)$, by (6.1) we have $(I^{\Gamma}-\mathcal{G}_{c}^{\Gamma}(s))(a,b)=\mathbbm{1}_{\{a\}}(b)-(\omega(b)-\mathbbm{1}_{\{\bar{a}\}}(b))^{-s}+\omega(b)^{-s}$. Hence,
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ such that $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and $\Lambda_{1}\cap\Lambda_{2}=\{c\}$, for some vertex $c\in V\Gamma$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
@@
The following lemma shows that the hypotheses of Proposition 6.4 are satisfied for every $c\in V\Gamma$ if, for instance, $\Gamma$ is a connected graph without $n$-cycles for every $n\geq 2$.
@@
Let $\{\Xi_{i}\}_{i\in I}$ be the collection of all the connected components of the graph $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})$. Recall that $\Gamma\setminus(o^{-1}(c)\cup\overline{o^{-1}(c)}\cup\{c\})=\bigsqcup_{i\in I}\Xi_{i}$. For every $i\in I$, there is exactly one edge $a_{i}\in o^{-1}(c)$ such that $t(a_{i})\in\Xi_{i}$. In fact, assume that there are $a,b\in o^{-1}(c)$ with $a\neq b$ such that $x=t(a),y=t(b)\in V\Xi_{i}$. Then the reduced path $[x,y]$ as in Remark 2.1.1 is contained in $\Xi_{i}$. Hence $a\cdot[x,y]\cdot\bar{b}$ is a cycle of length $\geq 2$ in $\Gamma$, impossible.
Consider subsets $\mathcal{E}_{1},\mathcal{E}_{2}$ of $o^{-1}(c)$ with the following properties: $\mathcal{E}_{1}\cap\mathcal{E}_{2}=\emptyset$, $o^{-1}(c)=\mathcal{E}_{1}\cup\mathcal{E}_{2}$ and, for every $k\in\{1,2\}$, every $1$-loop $a$ in $\mathcal{E}_{k}$ satisfies $\bar{a}\in\mathcal{E}_{k}$. Provided $\overline{\mathcal{E}}_{k}=\{\bar{a}\mid a\in\mathcal{E}_{k}\}$ for every $k\in\{1,2\}$, note that $(\mathcal{E}_{1}\cup\overline{\mathcal{E}}_{1})\cap(\mathcal{E}_{2}\cup\overline{\mathcal{E}}_{2})=\emptyset$. Moreover, if $|o^{-1}(c)|\geq 2$, take $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ so that $\mathcal{E}_{1}\neq\emptyset$ and $\mathcal{E}_{2}\neq\emptyset$. For $k\in\{1,2\}$, set also $I_{k}:=\{i\in I\mid a_{i}\in\mathcal{E}_{k}\}$. Note that $I_{1}\cap I_{2}=\emptyset$ and $I=I_{1}\cup I_{2}$. For $k\in\{1,2\}$, define the following subgraph of $\Gamma$:
$\Gamma_{k}:=\{c\}\cup\mathcal{E}_{k}\cup\overline{\mathcal{E}}_{k}\cup\bigsqcup_{i\in I_{k}}\Xi_{i}.$
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$]. Consider subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ satisfying $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ and such that $\Gamma_{3}:=\Gamma_{1}\cap\Gamma_{2}$ is a $1$-segment with edge set $\{a,\bar{a}\}$. Then,
@@
Note that $c\in V\Lambda_{1}$, $d\in V\Lambda_{2}$ and $V\Lambda_{1}\cap V\Lambda_{2}=\emptyset$. In particular, no edges of $\Lambda_{1}$ (resp. $\Lambda_{2}$) end in a vertex of $\Lambda_{2}$ (resp. $\Lambda_{1}$). We claim that
@@
From (6.18) (and (6.17)) we observe that $A(a,a)=1+\alpha^{-s}$ and $B(\bar{a},\bar{a})=1+\beta^{-s}$. Moreover, by (6.18) and (6.19), we have the following decompositions in diagonal blocks:
[…]
@@
(6.23) $\mathcal{Z}_{\Gamma_{1},a\to a}(s)^{-1}\hskip-2.84544pt=1-\beta^{-s}B(\bar{a},\bar{a})^{-1}-\alpha^{-s}A^{-1}(a,a)-\hskip-19.91684pt\sum_{b\in o^{-1}(c)\cap E\Lambda_{1}}\hskip-14.22636pt\mathcal{E}^{\Gamma_{1}}(s)(\bar{a},b)\cdot A^{-1}(b,a).$
@@
(6.24) $\mathcal{Z}_{\Gamma_{2},a\to a}(s)^{-1}\hskip-2.84544pt=1-\alpha^{-s}A(a,a)^{-1}-\beta^{-s}B^{-1}(\bar{a},\bar{a})-\hskip-19.91684pt\sum_{b\in o^{-1}(d)\cap E\Lambda_{2}}\hskip-14.22636pt\mathcal{E}^{\Gamma_{2}}(s)(a,b)\cdot B^{-1}(b,\bar{a}).$
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$]. Assume that there are subgraphs $\Lambda_{1}$ and $\Lambda_{2}$ of $\Gamma$ satisfying $\Gamma=\Lambda_{1}\cup\Lambda_{2}$ and such that $\Lambda_{1}\cap\Lambda_{2}$ is a $1$-segment graph with edge set $\{a,\bar{a}\}$. Then, for all subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$ such that $\Gamma_{i}\supseteq\Lambda_{i}$ for every $i\in\{1,2\}$, we have
@@
Let $\Gamma_{1}$ and $\Gamma_{2}$ be as in the statement. For simplicity, set $\Gamma_{3}=\Gamma_{1}\cap\Gamma_{2}$ and $\Gamma_{a}=\Lambda_{1}\cap\Lambda_{2}$. Since $\Lambda_{1}\subseteq\Gamma_{1}$ and $\Lambda_{2}\subseteq\Gamma_{2}$, we have $\Lambda_{1}\cap\Gamma_{2}=\Lambda_{1}\cap\Gamma_{3}$ and $\Lambda_{2}\cap\Gamma_{1}=\Lambda_{2}\cap\Gamma_{3}$. Therefore, the following decompositions hold:
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$]. Consider $a\in E\Gamma$ with $o(a)=:c\neq d:=t(a)$, and assume that $c$ is a terminal vertex in $\Gamma$. Put $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and denote by $\Lambda$ the graph obtained from $\Gamma$ by removing $a$, $\bar{a}$ and $c$. Then,
@@
Let $s\in\mathbb{C}$ be such that $\mathcal{Z}_{\Gamma,a\to a}(s)\neq 0$, and consider a total order $\leq$ on $E\Gamma$ such that $a<\bar{a}<b$ for all $b\in E\Lambda$. Then $\mathcal{E}^{\Gamma}(s)$ admits the following block decomposition:
@@
where $\underline{0}$ denotes the row zero vector in $\mathbb{C}^{|E\Lambda|}$. Moreover, denote by $E_{21}$ and $\underline{1}$ the elementary matrix associated to $(2,1)$ in $\mathrm{Mat}_{2}(\mathbb{C})$ and the matrix in $\mathrm{Mat}_{2}(\mathbb{C})$ with all entries equal to $1$, respectively. Hence,
[…]
@@
[…]
@@
(6.35) $\begin{split}\mathcal{Z}_{\Gamma,c\to c}(s)^{-1}&=\frac{\det(I^{\Gamma}-\mathcal{E}^{\Gamma}(s))}{\det(I-\mathcal{E}^{\Gamma}(s)+\mathcal{U}_{c,c}^{\Gamma}(s))}\\ &\hskip-19.91684pt=\frac{\det\Big(I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=:X}\Big)}{\det\Big(I_{2}-A+(\alpha+1)^{-s}E_{21}-\underbrace{B(I_{\Lambda}-D)^{-1}C}_{=X}\Big)}\\ &\hskip-19.91684pt=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1-X(a,a))(1-X(\bar{a},\bar{a}))+(\xi(\alpha,s)-X(\bar{a},a))(\beta^{-s}+X(a,\bar{a}))}.\end{split}$
@@
(6.36) $\begin{split}\mathcal{Z}_{\Gamma,a\to a}(s)^{-1}&=\frac{\det(I^{\Gamma}-D)}{\det(I-D+\mathcal{U}_{a,a}^{\Gamma}(s))}\\ &\hskip-19.91684pt=\frac{\det\Big(I_{2}-A-\overbrace{B(I_{\Lambda}-D)^{-1}C}^{=X}\Big)}{\det\Big(I_{2}+(\underline{1}-I_{2})A+\underbrace{(\underline{1}-I_{2})B(I_{\Lambda}-D)^{-1}C}_{=:Y}\Big)}\\ &\hskip-19.91684pt=\frac{(1-X(a,a))(1-X(\bar{a},\bar{a}))-(X(a,\bar{a})+\beta^{-s})(X(\bar{a},a)+\alpha^{-s})}{(1+\alpha^{-s}+Y(a,a))(1+\beta^{-s}+Y(\bar{a},\bar{a}))-Y(a,\bar{a})Y(\bar{a},a)}.\end{split}$
@@
Moreover, by Theorem 5.14, Fact 5.15 and since $\mathcal{E}^{\Lambda}(s)=D$,
$\begin{split}X(a,\bar{a})=Y(\bar{a},\bar{a})&=B_{1}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}C^{2}\\ &=(\beta+1)^{-s}\cdot\hskip-14.22636pt\sum_{b\in o^{-1}(d)\cap E\Lambda}\hskip-14.22636pt\omega(b)^{-s}e_{b}^{\Lambda}(I_{\Lambda}-\mathcal{E}^{\Lambda}(s))^{-1}(e_{d}^{\Lambda})^{t}\\ &=(\beta+1)^{-s}(\mathcal{Z}_{\Lambda,d\to d}(s)-1).\qed\end{split}$
@@
Let $(\Gamma,\omega)$ satisfy Setting $\Gamma$], and consider $a\in E\Gamma$ with $o(a)=t(a)=c$. Set $\omega(a)=\alpha+1$, $\omega(\bar{a})=\beta+1$, and consider the subgraph of $\Gamma$ given by $\Lambda:=\Gamma\setminus\{a,\bar{a}\}$. Then,
@@
[…]
@@
In this case $X=Y$, because $B_{1}=B_{2}$ and then $(\underline{1}-I_{2})B=B$. Moreover, recalling Theorem 5.14 and Fact 5.15, we have